Factor completely:
x^2+6x+9

Answer by malakumar_kos@yahoo.com(315)


x^2+6x+9 Factorise
x^2+6x+9 = x^2+3x+3x+9 = x(x+3)+3(x+3)
= (x+3).(x+3)
= (x+3)^2